OpenStudy (anonymous):

please help Rewrite 1−cos135 sin135 using a half-angle identity.

OpenStudy (anonymous):

Rewrite 1−cos135/sin135 using a half-angle identity.

OpenStudy (zehanz):

Do you know any formulas for half angles?

OpenStudy (zehanz):

BTW, your original question is different from the second

OpenStudy (anonymous):

cos^2 X= 1/2(1+cos(2X)) sin^2 x= 1/2(1-cos(2X))

OpenStudy (anonymous):

yeah it's the second one

OpenStudy (anonymous):

tan^2 X= 1-cos(2X)/1+cos(2X)

OpenStudy (zehanz):

This problem seems a bit far-fetched to me, because the exact values of this particular sin and cos are easily found. You might know them: \(\cos 135°=-\frac{1}{2}\sqrt{2}\), \(\sin 135°=\frac{1}{2}\sqrt{2}\), So the answer is 1 - - 1 = 2. I cannot see how half-angle formulas would be of help here...

OpenStudy (anonymous):

because it has to be re written

OpenStudy (anonymous):

thank you anyways

OpenStudy (zehanz):

OK, I will give it a try...(might take a few minutes)

OpenStudy (anonymous):

its ok no need to

OpenStudy (zehanz):

Sigh! (of relief)

OpenStudy (anonymous):

lol thank you :)